OpenStudy (anonymous):

what percent of 75 is 15?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

33%

OpenStudy (anonymous):

20 percent

OpenStudy (anonymous):

I have a question?

OpenStudy (anonymous):

im wrong dont listen to sorry

OpenStudy (johnweldon1993):

\[\large \frac{15}{75} = ?\]

OpenStudy (anonymous):

whats this fan thing it keeps saying I have 0

OpenStudy (anonymous):

People have to fan you

OpenStudy (anonymous):

you have 6 fans and i am one of them

OpenStudy (anonymous):

20%

OpenStudy (anonymous):

For you to get fans.

OpenStudy (ibbutibbu.):

First of all, the easiest thing for me to do is to turn this into a proportion. SO, 15/75 = x/100. The you multiply 15 and 100. The you divide the product by 75. Then you get your answer.

OpenStudy (uri):